# Ranking Large Language Models without Ground Truth

### v2

@@ Abstract
Evaluation and ranking of large language models (LLMs) has become an important problem with the proliferation of these models and their impact. Evaluation methods either require human responses which are expensive to acquire or use pairs of LLMs to evaluate each other which can be unreliable. In this paper, we provide a novel perspective where, given a dataset of prompts (viz. questions, instructions, etc.) and a set of LLMs, we rank them without access to any ground truth or reference responses. Inspired by real life where both an expert and a knowledgeable person can identify a novice our main idea is to consider triplets of models, where each one of them evaluates the other two, correctly identifying the worst model in the triplet with high probability. We also analyze our idea and provide sufficient conditions for it to succeed. Applying this idea repeatedly, we propose two methods to rank LLMs. In experiments on different generative tasks (summarization, multiple-choice, and dialog), our methods reliably recover close to true rankings without reference data. This points to a viable low-resource mechanism for practical use.
@@ 2 Related Work
Model Selection There are benchmarking deployments in the wild such as Chatbot Arena [^1] https://chat.lmsys.org/ Zheng et al. (2023) where humans provide the input prompts as well as provide their preference based on the output of two models. Chatbot Arena uses the Elo rating system Elo (1967), popularly used for rating Chess players, to rank LLMs. While collecting the input prompts to create a dataset is time-consuming, it is even more challenging to collect human labels, judgements, or preferences, particularly in cases where the deployment of models need to be made domain-specific. The number of applications and use cases may be so varied, and the human expertise need may be so high, that it will be too expensive to collect such data.
LLM-as-a-judge Trusted LLMs have been used as a proxy for human preferences and feedback Chiang and Lee (2023). Collecting human preferences are costly, however it has been shown metrics associated with benchmarks are not rich enough to capture the many and nuanced aspects that quantify what represents a good output a person might prefer. PandaLM Wang et al. (2023b) is trained on human-annotations in order to evaluate model performance on criteria difficult to capture in existing benchmarks reflecting human preferences in order to select a model out of several candidates. Other applications include acting as substitute for a traditional measurement of a user defined criteria. Instead of creating a traditional classification or regression model, the practitioner can specify the custom criteria in free text  [^2] https://python.langchain.com/docs/guides/evaluation/string/criteria_eval_chain. More generally, LLM-as-a-judge has also been used to rank model performance Chiang et al. (2023).
@@ 3.1 Greedy Triplet Ranking
Algorithm 1 Greedy Triplet Ranking (GTR) algorithm.
@@ 3.2 Full Triplet Ranking
Algorithm 2 Full Triplet Ranking (FTR) algorithm
@@ 4 Experiments
We test on two datasets: the CNN-Daily Mail (CNN/DM) summarization dataset Hermann et al. (2015); Nallapati et al. (2016) and XSUM Narayan et al. (2018). In both cases we gather responses from $40$ LLMs for $3000$ instances from HELM Liang et al. (2023). Our evaluation functions, i.e. $f$ in GTR and $g$ in FTR, employ ROUGE scores which are widely used for summarization tasks and in HELM. In particular, we consider the f-measure of bigrams. For the ‘most-common answer’, we use the top-$k$ bigrams from all model responses and consider this to be the reference answer ($k=256$ in our experiments). True ranking, against which our estimated ranking is compared, is inferred by comparing model outputs to the reference response in the dataset.
@@
We simulate responses from models with known accuracy in a multiple choice setting. Datasets are constructed for several cases by varying the number of models, model accuracy, number of prompts and number of possible answers. Since answers are discrete, the evaluation functions (i.e. $f$ in GTR and $g$ in FTR) use the equality operator to determine if a response is correct. We also experimented with a noisy equality operator, where the outcome of the evaluation is flipped with a known probability. This simulates a realistic case where comparisons are imperfect.
@@ 4.1 Results
Since the data in this case is simulated, we tested the impact of model performance on ranking recovery. Table 2 shows that even when the correctness of outcomes are pure chance (i.e. 50%) for the best of the models being ranked, the RBO metric is at $0.832$ for FTR and $0.723$ for GTR which indicate good agreement with true rankings. Moreover, as seen in Figure 13 in the appendix, FTR and even GTR are more robust than MCA at low-medium levels of noise. This suggests that our algorithms could still be used for non-generative or single token multiple choice kind of tasks.
@@
Our last experiment is qualitative and ground in a practical application of assessing amongst several fine-tuned variants. Specifically, from a pre-trained model ($PT$), several reward functions were specified that aimed to fine tune for multiple moral objectives. Three variants ($M_{1}$, $M_{2}$, and $M_{3}$) were LLMs created by fine-tuning for each reward function using RLHF. Outputs of each model for $100$ prompts were recorded and scored by human annotators. The annotator task was simply to choose the ‘best’ response, without qualifying any specific dimension for evaluation. Three annotators unrelated to this paper specified their preferences. We evaluated the FTR method using entailment probability and ROUGE scores as the basis for the evaluation function $g$. We ran only FTR here, since there were only four models and so efficiency wasn’t a factor. Our method ranked the models $[M_{1},M_{3},M_{2},PT]$ as compared to the human annotators $[M_{1},M_{2},M_{3},PT]$. Thus, we were able to rank the best and the worst models correctly. The middle two were quite close in performance from what we heard so flipping them was probably not that damaging.
@@ 5 Discussion
There are many interesting avenues to explore in the future. First, how we can incorporate and benefit from additional information such as a partial ordering between LLMs, which is reasonable to assume when some LLMs belong to a model family. The additional information could also be in the form of a few ground truth labels for the task at hand. In this case, there may be a wider set of possible methods such as those based on uncertainty or possibly even Bayesian approaches. Second, rather than triplets what if we consider larger sets to compare in each round. Of course, for FTR the number of comparisons might scale exponentially with the set size so more efficient variants might have to be thought of. Here it is possible that the variance in ranking the worse models might increase as a larger set of judges will produce their own ranking. Moreover, asking each judge to rank a larger set of models might be more error prone. Nonetheless, the bias-variance trade-off between different set sizes might be interesting to explore. Third, more efficient variants than GTR may be worth exploring. For instance, in GTR the top two models at the end of the first round may be used to rank the rest resulting in a linear time variant. Also methods from matrix completion could be explored where we obtain a few partial orderings running GTR or FTR for a few iterations. Finally, our approach might find use in other contexts where, consolidation of responses into a single response is not easy such as when evaluating non-expert human annotators on descriptive tasks, or even in accelerated drug discovery type of applications where different algorithms might recommend different molecular structures that may be hard to consolidate.
@@ A.1 Summarization
Figure 10 shows sample bump charts that visualize rank changes between the ‘true’ (based on reference data) and the estimated ranks from the proposed methods. Figure 9(a) shows perfect agreement between estimated and true rankings. In Figure 9(b) rankings are not exactly recovered, but the clusters of good and bad models are persisted, even if the exact rankings of them within the clusters are not correct. This is also seen in Figure 9(c) and 9(d).

### v3

@@ Abstract
Evaluation and ranking of large language models (LLMs) has become an important problem with the proliferation of these models and their impact. Evaluation methods either require human responses which are expensive to acquire or use pairs of LLMs to evaluate each other which can be unreliable. In this paper, we provide a novel perspective where, given a dataset of prompts (viz. questions, instructions, etc.) and a set of LLMs, we rank them without access to any ground truth or reference responses. Inspired by real life where both an expert and a knowledgeable person can identify a novice our main idea is to consider triplets of models, where each one of them evaluates the other two, correctly identifying the worst model in the triplet with high probability. We also analyze our idea and provide sufficient conditions for it to succeed. Applying this idea repeatedly, we propose two methods to rank LLMs. In experiments on different generative tasks (summarization, multiple-choice, and dialog), our methods reliably recover close to true rankings without reference data. This points to a viable low-resource mechanism for practical use [^1] Code available at https://huggingface.co/spaces/ibm/llm-rank-themselves..
@@ 2 Related Work
Model Selection
There are benchmarking deployments in the wild such as Chatbot Arena [^2] https://chat.lmsys.org/ Zheng et al. (2023) where humans provide the input prompts as well as provide their preference based on the output of two models. Chatbot Arena uses the Elo rating system Elo (1967), popularly used for rating Chess players, to rank LLMs. While collecting the input prompts to create a dataset is time-consuming, it is even more challenging to collect human labels, judgements, or preferences, particularly in cases where the deployment of models need to be made domain-specific. The number of applications and use cases may be so varied, and the human expertise need may be so high, that it will be too expensive to collect such data.
LLM-as-a-judge
Trusted LLMs have been used as a proxy for human preferences and feedback Chiang and Lee (2023). Collecting human preferences are costly, however it has been shown metrics associated with benchmarks are not rich enough to capture the many and nuanced aspects that quantify what represents a good output a person might prefer. PandaLM Wang et al. (2023b) is trained on human-annotations in order to evaluate model performance on criteria difficult to capture in existing benchmarks reflecting human preferences in order to select a model out of several candidates. Other applications include acting as substitute for a traditional measurement of a user defined criteria. Instead of creating a traditional classification or regression model, the practitioner can specify the custom criteria in free text  [^3] https://python.langchain.com/docs/guides/evaluation/string/criteria_eval_chain. More generally, LLM-as-a-judge has also been used to rank model performance Chiang et al. (2023).
@@ 3.1 Greedy Triplet Ranking
       $T=\{M_{1},M_{2},M_{3}\}$ Triplet to begin for $i=3$ to $|M|$ do
             $T\leftarrow T\cup M_{i}$ $W\leftarrow$ Using $f(T,D)$ compare pairs of models with the third being the judge and output the model voted as the worst by the other two. In case of tie output $M_{i}$. $T\leftarrow T\setminus W$
      $M\leftarrow M\setminus T$ Remove top two models as they will be added to $R$ if $R==\phi$ then
            $R\leftarrow list(T)$ Amongst top two models randomly pick one as the best.
      else
             $T\leftarrow T\cup R_{1}$ $W\leftarrow$ Using $f(T,D)$ and $R_{1}$ as judge output the worse model. $R\leftarrow$ append $(T\setminus W,W)$ to $R$
if $|M|==1$ then
       $R\leftarrow$ append $M$ to $R$
else if $|M|==2$ then
      $T\leftarrow M\cup R_{1}$ $W\leftarrow$ Using $f(T,D)$ and $R_{1}$ as judge output the worse model. $R\leftarrow$ append $(T\setminus W,W)$ to $R$
Output: $R$
Algorithm 1 Greedy Triplet Ranking (GTR).
@@ 3.2 Full Triplet Ranking
       $m_{ij}=\frac{1}{|M|}\sum_{k}y_{ijk}r_{k}\,\forall M_{i},M_{j}\in M$ Weighted preference matrix. $z_{ij}=1$ if $m_{ij}\geq m_{ji}\,\forall M_{i},M_{j}\in M$ and 00 otherwise # Majority vote. $r^{\prime}_{i}=\frac{1}{|M|-1}\sum_{j}z_{ij}\,\forall M_{i}\in M$ New reputation score. $\delta=\sum_{k}|r_{k}-r^{\prime}_{k}|$ Convergence condition. if $\delta\leq\epsilon$ then
             Break
      else
             $r_{k}=r^{\prime}_{k}$ Update reputation
Sort $M$ using reputation score $r_{k}$ for ranking $R$. Output: $R$
Algorithm 2 Full Triplet Ranking (FTR).
@@ 4 Experiments
We test on two datasets: the CNN-Daily Mail (CNN/DM) summarization dataset Hermann et al. (2015); Nallapati et al. (2016) and XSUM Narayan et al. (2018). In both cases we gather responses from $40$ LLMs for $3000$ instances from HELM Liang et al. (2023). Our evaluation functions, i.e. $f$ in GTR and $g$ in FTR, employ ROUGE scores which are widely used for summarization tasks and in HELM. In particular, we consider the f-measure of bigrams. For the ‘most-common answer’, we use the top-$k$ bigrams from all model responses and consider this to be the reference answer ($k=256$ in our experiments). For this case, we also compare with LLM-as-a-judge approach using Prometheus-2 Kim et al. (2024), which is a state-of-the-art LLM evaluator or judge model. True ranking, against which our estimated ranking is compared, is inferred by comparing model outputs to the reference response in the dataset.
@@
We simulate responses from models with known accuracy in a multiple choice setting. Datasets are constructed for several cases by varying the number of models, model accuracy, number of prompts and number of possible answers. Table 5 shows a sample. Since answers are discrete, the evaluation functions (i.e. $f$ in GTR and $g$ in FTR) use the equality operator to determine if a response is correct. We also experimented with a noisy equality operator, where the outcome of the evaluation is flipped with a known probability. This simulates a realistic case where comparisons are imperfect.
@@ 4.1 Results
Since the data in this case is simulated, we tested the impact of model performance on ranking recovery. Table 2 shows that even when the correctness of outcomes are pure chance (i.e. 50%) for the best of the models being ranked, the RBO metric is at $0.832$ for FTR and $0.723$ for GTR which indicate good agreement with true rankings.
@@
We investigate the role of noise in the evaluation function. In multiple choice, this comparison is exact as it uses an equality operator. This is substituted with a noisy equality operator, which randomly flips the outcome with a noise probability $p$. This allows us to evaluate the robustness of our methods in the face of noisy evaluations - which is a more realistic setting in generative cases. Figure 4 shows the RBO metric as a function of noise for various methods. Both FTR and GTR are much more robust to noise at low and medium levels as compared to MCA. FTR does best here with its dynamic weighted voting making it less susceptible to noise. MCA on the other hand starts to degrade rapidly as noise increases. Additional plots are presented in Figure 14 in the appendix. This suggests our algorithms have broader applicability, including for single token classification type tasks.
Figure 4: Quality of rankings recovered as a function of noise in the evaluation function for different methods. Four sets of $10$ LLMs with each set having a maximum LLM accuracy of $30\%$, $50\%$, $70\%$ or $90\%$ are considered, where the number of questions is $50$. We see the robustness of the proposed methods (GTR and FTR) at low to medium levels of noise in such a setup.
@@
Our last use case is qualitative and ground in a practical application of assessing amongst several fine-tuned variants. Specifically, from a pre-trained model ($PT$), several reward functions were specified that aimed to fine tune for multiple moral objectives. Three variants ($M_{1}$, $M_{2}$, and $M_{3}$) were LLMs created by fine-tuning for each reward function using RLHF. Outputs of each model for $100$ prompts were recorded and scored by human annotators. The annotator task was simply to choose the ‘best’ response, without qualifying any specific dimension for evaluation. Three annotators unrelated to this paper specified their preferences. We evaluated the FTR method using entailment probability and ROUGE scores as the basis for the evaluation function $g$. We ran only FTR here, since there were only four models and so efficiency wasn’t a factor. Our method ranked the models $[M_{1},M_{3},M_{2},PT]$ as compared to the human annotators $[M_{1},M_{2},M_{3},PT]$. Thus, we were able to rank the best and the worst models correctly. The middle two were quite close in performance from what we heard so flipping them was probably not that damaging.
Comparison with LLM-as-a-judge
Next, we compare our approach to the growing LLM-as-a-judge literature Zheng et al. (2024). In this setting, language models themselves are used as evaluators. LLM judges used in this was offer lower cost of annotations/labels compared to human annotations and can make metric-free evaluations, for instance evaluating responses that meet criteria such as ‘child-friendliness’ or ‘professional tone’. Prometheus Kim et al. (2023) and its subsequent version Kim et al. (2024) are two such specialized LLM evaluators.
We leverage Prometheus-2 to compare LLM outputs in a pairwise fashion. As such, we prompt it to compare the results of two summaries from two LLMs for the CNN/DM summarization dataset. This is done for all pairs of LLMs and for a sample of $50$ questions in the benchmark. For each evaluation, the LLM judge (i.e. Prometheus-2) declares a winning LLM. We aggregate win rates across the benchmark and derive a LLM ranking based on win rates. We consider a (random) sample of $10$ LLMs from the benchmark. Our task criteria is simply ‘Which is the better response?’.
Table 3 shows the mean and standard deviation of ranking metrics compared to this approach. Our methods out-perform the LLM judge across all metrics. Additional details are included in the Appendix C.
Table 3: Comparison with state-of-the-art Prometheus-2 LLM-judge using pairwise evaluation on CNN/DM summarization task with $10$ models. Best (mean) result in bold.
Algorithm RBO MAP-3 MAP-5
FTR 0.835$\pm$0.04 0.200$\pm$0.14 0.523$\pm$0.16
GTR 0.841$\pm$0.06 0.300$\pm$0.26 0.448$\pm$0.23
Prometheus-2 0.806$\pm$0.07 0.200$\pm$0.18 0.354$\pm$0.28
MCA 0.792$\pm$0.05 0.144$\pm$0.16 0.296$\pm$0.26
@@ 5 Discussion
There are many interesting avenues to explore in the future. First, how we can incorporate and benefit from additional information such as a partial ordering between LLMs, which is reasonable to assume when some LLMs belong to a model family. The additional information could also be in the form of a few ground truth labels for the task at hand. In this case, there may be a wider set of possible methods such as those based on uncertainty or possibly even Bayesian approaches. Second, rather than triplets what if we consider larger sets to compare in each round. Of course, for FTR the number of comparisons might scale exponentially with the set size so more efficient variants might have to be thought of. Here it is possible that the variance in ranking the worse models might increase as a larger set of judges will produce their own ranking. Moreover, asking each judge to rank a larger set of models might be more error prone. Nonetheless, the bias-variance trade-off between different set sizes might be interesting to explore. Third, more efficient variants than GTR may be worth exploring. For instance, in GTR the top two models at the end of the first round may be used to rank the rest resulting in a linear time variant. Also methods from matrix completion could be explored where we obtain a few partial orderings running GTR or FTR for a few iterations. Finally, our approach might find use in other contexts where, consolidation of responses into a single response is not easy such as: i) when evaluating non-expert human annotators on descriptive tasks. ii) In accelerated drug discovery type of applications where different algorithms might recommend different molecular structures that may be hard to consolidate or iii) even in multi-agent systems where a protocol is needed to decide which agent to follow in a specific context.
@@ A.1 Summarization
All results in the main paper are reported based on the random sampling strategy. Figures 9 and 10 show the assessment when the sampling strategy changes.
@@
Figure 11 shows sample bump charts that visualize rank changes between the ‘true’ (based on reference data) and the estimated ranks from the proposed methods. Figure 11(a) shows perfect agreement between estimated and true rankings. In Figure 11(b) rankings are not exactly recovered, but the clusters of good and bad models are persisted, even if the exact rankings of them within the clusters are not correct. This is also seen in Figure 11(c) and 11(d).
@@ A.2 Multiple Choice
A.3 Models used
Table 4 shows the list of models used in the HELM experiments for summarization.
No. Model
1 AlephAlpha_luminous-base
2 AlephAlpha_luminous-extended
3 AlephAlpha_luminous-supreme
4 ai21_j1-grande
5 ai21_j1-grande-v2-beta
6 ai21_j1-jumbo
7 ai21_j1-large
8 ai21_j2-grande
9 ai21_j2-jumbo
10 ai21_j2-large
11 anthropic_stanford-online-all-v4-s3
12 cohere_command-medium-beta
13 cohere_command-xlarge-beta
14 cohere_large-20220720
15 cohere_medium-20220720
16 cohere_medium-20221108
17 cohere_small-20220720
18 cohere_xlarge-20220609
19 cohere_xlarge-20221108
20 microsoft_TNLGv2_530B
21 microsoft_TNLGv2_7B
22 openai_ada
23 openai_babbage
24 openai_curie
25 openai_davinci
26 openai_gpt-3.5-turbo-0301
27 openai_text-ada-001
28 openai_text-babbage-001
29 openai_text-curie-001
30 openai_text-davinci-002
31 openai_text-davinci-003
32 together_bloom
33 together_gpt-j-6b
34 together_gpt-neox-20b
35 together_opt-175b
36 together_opt-66b
37 together_redpajama-incite-base-3b-v1
38 together_yalm
39 writer_palmyra-instruct-30
40 writer_palmyra-x
Table 4: Models used in summarization experiments
A.4 Synthetic Data Generation
Table 5 shows an example of synthetic data generated for the multiple choice setting. We parameterize the generation on model accuracy, number of possible answers, and the number of questions in the dataset.
GT M1 M2 M3 M4 M5
Q0 0 0 0 0 7 0
Q1 2 6 2 2 2 1
Q2 1 1 1 1 0 7
Q3 6 6 6 4 6 6
Q4 5 5 1 5 5 5
Table 5: Sample generated dataset with model accuracies in the range $50-80\%$, with $10$ possible answers per question. GT denotes the ground truth, i.e. true label, and M1-5 denote simulated model responses.
Appendix B Most Common Answer
We now present additional details on the Most-Common Answer (MCA) method that we use as a baseline. When the context is multiple choice, the concept of the most common answer is straight forward. You assign the label with the majority vote from all the LLMs being tested as the true label. For instance, in a multiple choice setting with $4$ options $A,B,C,D$, if from five different models you gather responses for a specific question as $A,A,B,C,D$, then $A$ would be considered as the true label.
In generative settings, such as for summarization, what constitutes a ‘common’ answer involves more work. To establish an MCA baseline, we first generate all n-grams from all model responses. We then use the most frequent n-grams across all responses upto a specific size, and consider that to be the ‘true’ response. Specific answers from the models are then evaluated against this to generate the closest response.
This is illustrated using a simple example. Consider three models’ responses to the question What is the capital of Canada? in Table 6. From all the responses, we compute frequencies of all bigrams to get ’ta’: 3, ’nt’: 2, ’Ot’: 2, ’tt’: 2, ’aw’: 2, ’wa’: 2,’To’: 1,’or’: 1,’ro’: 1,’on’: 1,’to’: 1,’a,’: 1,’, ’: 1,’ O’: 1,’On’: 1,’ar’: 1,’ri’: 1,’io’: 1. From this, we select the top-k bigrams to constitute the most-common answer. In this example we assume the top five, but in the summarization experiments we consider k to be 256. We can now compute similarity measures of each model response relative to this most-common answer. Rouge-2 scores, as an example are shown in Table 6 indicating that M3 has the best response.
Model Response Rouge-2 w.r.t. MCA
M1 Toronto 0.117
M2 Ottawa, Ontario 0.480
M3 Ottawa 0.500
MCA ta: 3, nt: 2, ot: 2,
tt: 2, aw: 2 -
Table 6: Illustration of model responses and most-common answer (MCA) for the question: ‘What is the capital of Canada?’
Appendix C LLM as a Judge
We prompt Prometheus-2 Kim et al. (2024) to compare the results of two summaries from the CNN/DM summarization dataset. This is done for all pairs of LLMs and for a sample of $50$ questions in the benchmark. For each evaluation, the LLM Judge declares a winning LLM. Our task criteria is simply ‘Which is the better response?’. Generally, human annotators judging summarization quality do so along specific attributes such as coherence, fluency, consistency, relevance, and Fabbri et al. (2021). We additionally compared Prometheus with human annotations along these dimensions and find poor agreement. For this task, the LLM Judge is overly pessimistic across all dimensions, see Figure 15.
Figure 15: Comparison of Human annotations with LLM Judge (Prometheus) annotations based on the SummEval dataset Fabbri et al. (2021).
Next we considered pairwise evaluations, specifically, we prompted Prometheus-2 with the following prompt template:
"""###Task Description:
An instruction (might include an Input inside it),
a response to evaluate, and a score rubric
representing  a evaluation criteria are given.
1. Write a detailed feedback that assess the
quality of two responses strictly based on the
given score rubric, not evaluating in general.
2. After writing a feedback, choose a better
response between Response 1 and Response 2.
You should refer to the score rubric.
3. The output format should look as follows:
"Feedback: (write a feedback for criteria)
[RESULT] (1 or 2)"
4. Please do not generate any other opening,
closing, and explanations.
###Instruction:
{instruction}
###Response 1:
{response_1}
###Response 2:
{response_2}
###Score Rubric:
{rubric}
###Feedback:
No additional checks for positional bias were performed.